The world's fastest humans can reach speeds of about 11 m/s. In order to increase his gravitational potential energy by an amount equal to his kinetic energy at full speed, how high would such a sprinter need to climb?

How do you set this up?
Answer #1
(1/2) m (121) = m (9.8) h

h = 121 /(9.8*2) = 6.17 meters
answered by: jordie
Answer #2
That really helped thank you very much
answered by: karn